March 26th 2021

From Hall A Wiki
Jump to: navigation, search
  • Updates
    • HCAL
      • cable GUI installed on enpcamsonne, section 7.5 on wiki for documentation
      • looking at cable length attenuation


      • gathering spool of cables and lemos to give to Wendy for 60 more cables for overlapping summing
      • Chris agreed to make a patch panel for the DAC
      • looking for the FADC being repaired
      • maybe look into setting up external clock for F1
    • Bigbite shower
      • Issue with Fastbus getting stuck in weird state. Bryan has a temporary fix.
      • 12 FADCs added, jumpers checked to be set to 2V and will start cabling
      • Provokar more data for preshower calibration and will switch to FADC
      • noise on one side of preshower, HV crate was replaced ( goes away when moving to other HV crate ) but still shifting pedestals on some channels, still need to look at signals with scope ( usually 5 mins runs ), might be ground loop - HV supposed to have same ground but need to check. Some channels have the shift on some channels with scope directly on PMT ( grounding near PMT did not help ), but disappear R26 when HV swapped to another HV card.

Schematics of bases of PMTs ? From Hermes not sure we have. Look with scope on HV too, see in Fastbus even with filter ( so higher frequency noise ).

    • Bigbite hodoscope
      • fix issues with analysis script - everything working
      • more data taking
      • Steve working on the TDC scaler using the hodoscope DAQ
    • GRINCH
      • still working on integration in SBS-offline , some work on best structure
      • software running but no data - need to check map
      • Juan Carlos has information for podd debugging to display data
      • already checked with datadecode that raw data is fine, so still need to fix decoding software
    • LHRS
      • problem with HV tripping on VDC, added more oil on bubbler fixed it
      • Maria working on the DAQ
      • eventually will port the FADC setup from Bigbite
    • CDet
      • issue with sbs2 computer but it was because UPS dead, now plugged directly so working now
      • taking data , students looking at data
      • Ralph set up CDet channel on slack
    • TEDF GEM
      • slides from
      • clock : daisy chaining from input to output - going through PLL and delay - Ben prefer star configuration to daisy chaining - input 50 ohm terminated low level TTL
      • just run on internal clock for the phase study
    • UVA GEM
      • APV front end and backplane being repaired by electronics group
      • tracking APV baseline issue - might be from 5 slot backplane - will try next week
    • TS crate cabled, need to switch to CODA3 Centos 7 and continue next week